### v1

@@ Abstract
This paper introduces SpoofCeleb, a dataset designed for Speech Deepfake Detection (SDD) and Spoofing-robust Automatic Speaker Verification (SASV), utilizing source data from real-world conditions and spoofing attacks generated by Text-To-Speech (TTS) systems also trained on the same real-world data. Robust recognition systems require speech data recorded in varied acoustic environments with different levels of noise to be trained. However, existing datasets typically include clean, high-quality recordings (bona fide data) due to the requirements for TTS training; studio-quality or well-recorded read speech is typically necessary to train TTS models. Existing SDD datasets also have limited usefulness for training SASV models due to insufficient speaker diversity. We present SpoofCeleb, which leverages a fully automated pipeline that processes the VoxCeleb1 dataset, transforming it into a suitable form for TTS training. We subsequently train 23 contemporary TTS systems. The resulting SpoofCeleb dataset comprises over 2.5 million utterances from 1,251 unique speakers, collected under natural, real-world conditions. The dataset includes carefully partitioned training, validation, and evaluation sets with well-controlled experimental protocols. We provide baseline results for both SDD and SASV tasks. All data, protocols, and baselines are publicly available at https://jungjee.github.io/spoofceleb.
@@ 1 INTRODUCTION
The quality of synthetic speech has improved rapidly in recent years, driven by advancements in technologies such as flow matching, neural codecs, and speech-language modeling [1, 2, 3]. These innovations have significantly enhanced the naturalness and intelligibility of generated speech. Furthermore, the increasing availability of open sources and APIs for Text-To-Speech (TTS) systems has made high-quality synthetic speech more accessible to the general public [4, 5].
@@
Although originally developed for positive applications, this technology is increasingly being exploited for malicious purposes [6, 7]. Synthetic speech generated with harmful intent, often referred to as spoofing, is being used to deceive individuals in scenarios such as voice phishing (or vishing). Additionally, spoofing undermines the reliability of speech biometric systems, including Automatic Speaker Verification (ASV), many of which remain highly vulnerable to such attacks [8, 9].
@@
In response to these challenges, several datasets have been developed to advance research in Speech Deepfake Detection (SDD) [10, 11, 12]. For robust recognition systems, it is essential to have training data that covers a wide range of real-world acoustic environments and speaker diversity. However, speech generation systems like TTS and Voice Conversion (VC) typically require studio-quality or clean, read speech for training. As a result, existing datasets tend to feature clean, monotonic bona fide speech, with spoofed samples also being clean, as they are synthesized using TTS and VC systems trained on such data. Additionally, the emerging task of Spoofing-robust Automatic Speaker Verification (SASV) [13] lacks dedicated datasets. Many SDD datasets also suffer from limited speaker diversity, which hinders research on SASV systems that require training with data from hundreds or even thousands of speakers.
@@
To this end, we introduce the SpoofCeleb dataset, which is built upon VoxCeleb1 [14], a widely-used Automatic Speaker Verification (ASV) dataset consisting of the voices of 1,25112511,2511 , 251 celebrities recorded in real-world conditions. We develop a fully automated pipeline that processes VoxCeleb1 to produce in-the-wild bona fide speech samples that can be used for training TTS systems. [^1] The development of this pipeline is extensive, and the resulting bona fide speech data can serve other purposes, such as advancing research on TTS systems trained on noisy, in-the-wild data. We detail this aspect in a separate work, referring to the dataset as TTS In The Wild (TITW) [15]. From the two available TTS training sets in TITW, we use TITW-Easy as the source dataset to generate 23 spoofing attacks. SpoofCeleb is the first dataset explicitly designed for both SDD and SASV, where the bona fide speech is real-world, noisy speech. The dataset is divided into three partitions for training, validation, and evaluation, accompanied by evaluation protocols. Baseline systems trained on SpoofCeleb’s training partition are also presented, demonstrating the dataset’s effectiveness and potential for future research in SDD and SASV.
@@
Table 1: List of datasets in Speech Deepfake Detection (SDD) and Spoofing-robust Automatic Speaker Verification (SASV). FakeAVCeleb [16] and “In The Wild [17]” also have in-the-wild data. However, they have either only an evaluation partition or the number of speakers or spoofing attacks is limited.
@@ 2 RELATED WORKS
Datasets for SDD and the generation-recognition trade-off. To safeguard the authenticity of speech, several corpora have been published to support research in SDD [10, 11, 18, 9, 26, 21, 12, 23]. One of the most critical decisions when creating these datasets is the selection of the source data (i.e., bonafide speech). This decision involves a trade-off, which we refer to as the “generation-recognition trade-off.”
@@
On the recognition side, for both SDD and SASV, incorporating data with diverse noise, reverberation, and varied domains is essential for training robust models. It is well known that recognition models trained solely on clean speech often struggle to generalize effectively to noisy environments during inference [18]. While data augmentation techniques can help mitigate this issue [31], the most effective solution is to use training data drawn from a wide range of real-world sources.
@@
Conversely, traditional TTS training requires a carefully curated and recorded dataset. Sentence prompts must be selected to ensure comprehensive phonetic coverage [32], and recordings are typically made by voice talents in clean environments, ideally in a single anechoic studio. These recordings are of high studio quality and are carefully articulated, but they are not scalable. For instance, the well-known CMU Arctic database includes recordings from fewer than 10 voice talents, each reading approximately 1,000 speech prompts [32]. Modern TTS systems, however, often require significantly more training data. Instead of relying on these small-scale, TTS-specific databases, contemporary models frequently use audiobook datasets (e.g., MLS [33]), which, while not studio-grade, consist of relatively clean audiobook recordings made by numerous readers in their homes or offices.
@@
Existing SDD datasets tend to lean towards the generation side of the generation-recognition trade-off. They use source datasets that consist of either studio-quality or high-quality speech, facilitating the training of TTS and Voice Conversion (VC) systems and the successful generation of spoofed speech samples. However, both the bonafide and spoofed speech in these datasets are exceedingly clean, making them far from real-world, noisy speech data.
@@
SpoofCeleb is the first dataset to utilize real-world, noisy, and reverberant data from VoxCeleb1 as the source for training and synthesizing spoofed speech. We tackle the generation-recognition trade-off by employing a carefully curated, fully automated pre-processing pipeline that allows TTS models to be trained on data that more closely mirrors real-world conditions.
@@
Datasets for SASV. As SASV is an emerging task extending the scope of ASV systems with spoofing robustness, dedicated datasets for SASV lack. Earlier works in SASV have relied on SDD datasets [34, 35]. However, existing SDD datasets do not prioritize speaker diversity and balance, both of which are critical for SASV. Additionally, most datasets lack a sufficient number of speakers.
@@
To the best of our knowledge, VSASV [30], a parallel effort to SpoofCeleb, is the only existing work attempting to address these limitations by creating a dataset specifically for SASV. SpoofCeleb complements VSASV while also offering several distinctions. While VSASV includes three spoofing attacks, SpoofCeleb contains 23. Although VSASV uses in-the-wild bonafide data, its spoofed data is derived from high-quality sources due to the challenges in developing TTS systems with in-the-wild data. In contrast, SpoofCeleb adopts VoxCeleb1, a widely-used ASV dataset recorded in the wild, as its bonafide source. Additionally, VSASV includes approximately 300k samples, whereas SpoofCeleb offers over 2.5 million samples. Table 1 compares SpoofCeleb with other datasets in the SDD and SASV fields.
@@
Figure 1: Overall process pipeline of SpoofCeleb dataset collection. (a): proposed fully automated pipeline transcribes, segments, filters, enhances, and again filters with DNSMOS to derive TITW-Easy [15] from VoxCeleb1 [14], adequate to be used for TTS training. (b): 23 different TTS systems are trained using TITW-Easy and spoof speech samples are generated. All generated spoofing samples are combined with TITW-Easy to constitute the SpoofCeleb dataset.
@@ 3 SOURCE DATASET: TITW
Our goal is to create a dataset for SDD and SASV using VoxCeleb1 as the source so that both bonafide and spoofed samples reflect real-world scenarios. However, VoxCeleb1 is not suitable for direct use in TTS training. [^2] Our preliminary attempts to train TTS systems using the raw VoxCeleb1 data without further processing were unsuccessful. The challenges with VoxCeleb1 are multifaceted; for example, the speech samples often (i) contain overly emotional expressions, (ii) include extended non-speech segments, or (iii) have excessively long durations. To address these issues, we developed a fully automated pipeline that processes VoxCeleb1 into the Text-To-Speech In The Wild (TITW) dataset, which can be used for TTS training.
Figure 1-(a) illustrates the automated processing pipeline we developed to generate the TITW dataset. The pipeline begins by transcribing and obtaining word-level alignment using the WhisperX toolkit [36]. This toolkit transcribes the speech using the pre-trained Whisper Large v2 Automatic Speech Recognition (ASR) model [37], while word-level segmentation is derived from another phoneme-based ASR model. For a small subset of randomly selected samples, we also transcribed the text using the OWSMv3.1 model [38] and cross-checked the accuracy of the transcriptions. We then segmented the utterances from VoxCeleb1 whenever a silence longer than 500ms was detected, resulting in multiple segments from a single utterance. Next, we applied a series of heuristic-driven rules – developed through several iterations of TTS training – to filter the data. We discarded any samples that (i) were non-English, (ii) were shorter than 1 second or longer than 8 seconds, (iii) contained one or more words with a duration exceeding 500ms, or (iv) had empty transcriptions.
After completing the initial processing steps (referred to as TITW-Hard in [15]), we conducted multiple iterations of TTS training trials. Despite these efforts, training remained extremely challenging for most TTS systems, with only a few recent models showing success. The generated speech was still insufficient to deceive pre-trained ASV systems, as measured by the SPooF Equal Error Rate (SPF-EER) metric [13]. [^3] The SPF-EER is calculated by assessing an ASV system’s ability to correctly accept target trials while rejecting spoofed non-target trials. Bona fide non-target trials are excluded from this protocol, as the focus is solely on evaluating the ASV system’s spoofing robustness. To address this, we applied speech enhancement using a pre-trained model, DEMUCS, and excluded samples with DNSMOS “BAK” scores below 3.0. The final number of speech segments (TITW-Easy in [15]) is approximately 248k, which serves as the bona fide portion of the SpoofCeleb dataset. For full details on the preparation of TITW from VoxCeleb1, refer to [15].
@@ 4 SPOOFCELEB
Figure 1-(b) illustrates the composition of the SpoofCeleb dataset. The TITW dataset serves as the foundation for training multiple TTS systems. These systems are then used to synthesize spoofed speech samples, which are combined with the bona fide speech samples from TITW to form the complete SpoofCeleb dataset. To achieve this, we utilize 4 acoustic models, 6 waveform models (i.e., vocoders), and 5 End-to-End (E2E) models. Unless mentioned otherwise, all models are trained from scratch using the TITW-Easy data.
@@ 4.1 Acoustic models
Training acoustic models using in-the-wild data was one of the most challenging aspects of the SpoofCeleb dataset collection. We applied several criteria to evaluate the success of the training, including (but not limited to) speech intelligibility, measured by Word Error Rate (WER), noisiness, assessed by DNSMOS, and speaker identity, evaluated using SPF-EER. Among these metrics, SPF-EER was prioritized as the primary measure, since the most critical factor in a spoofing attack is whether it can deceive an ASV system. The final models that were successfully trained include TransformerTTS, GradTTS, Matcha-TTS, and BVAE-TTS.
TransformerTTS  [39] is an autoregressive TTS model that generates mel-spectrograms from textual input using a transformer-based architecture. The model employs a sequence of transformer encoder and decoder blocks with multi-head self-attention. We trained TransformerTTS using the ESPnet toolkit [40]. [^4] https://github.com/ESPnet/ESPnet.
GradTTS.  [41] is a TTS model with a score-based decoder that generates mel-spectrograms by gradually transforming noise predicted by the text encoder. During inference, we set the denoise step to 50 to ensure high-quality speech generation. We utilize the official implementation and follow the default settings. [^5] https://github.com/huawei-noah/Speech-Backbones.
Matcha-TTS. [42] is an efficient non-autoregressive TTS model based on an optimal-transport conditional flow matching decoder [1]. Unlike score-based models, it constructs a more direct sampling trajectory, allowing for high-quality generation with fewer sampling steps. We utilized the official implementation. [^6] https://github.com/shivammehta25/Matcha-TTS.
BVAE-TTS.  [43] utilizes a Bidirectional-inference Variational AutoEncoder (BVAE) to model the hierarchical relationships between text and speech. By leveraging the attention maps generated by BVAE-TTS, the model jointly trains a duration predictor, enabling robust and efficient non-autoregressive speech generation. We utilized the official implementation. [^7] https://github.com/LEEYOONHYUNG/BVAE-TTS.

### v2

@@ Abstract
This paper introduces SpoofCeleb, a dataset designed for Speech Deepfake Detection (SDD) and Spoofing-robust Automatic Speaker Verification (SASV), utilizing source data from real-world conditions and spoofing attacks generated by Text-To-Speech (TTS) systems also trained on the same real-world data. Robust recognition systems require speech data recorded in varied acoustic environments with different levels of noise to be trained. However, current datasets typically include clean, high-quality recordings (bona fide data) due to the requirements for TTS training; studio-quality or well-recorded read speech is typically necessary to train TTS models. Current SDD datasets also have limited usefulness for training SASV models due to insufficient speaker diversity. SpoofCeleb leverages a fully automated pipeline we developed that processes the VoxCeleb1 dataset, transforming it into a suitable form for TTS training. We subsequently train 23 contemporary TTS systems. SpoofCeleb comprises over 2.5 million utterances from 1,251 unique speakers, collected under natural, real-world conditions. The dataset includes carefully partitioned training, validation, and evaluation sets with well-controlled experimental protocols. We present the baseline results for both SDD and SASV tasks. All data, protocols, and baselines are publicly available at https://jungjee.github.io/spoofceleb.
@@ 1 INTRODUCTION
The quality of synthetic speech has improved rapidly, driven by advancements in technologies such as flow matching, neural codecs, and speech-language modeling [1, 2, 3]. These innovations have significantly enhanced the naturalness and intelligibility of generated speech. The increasing availability of open sources and APIs for Text-To-Speech (TTS) systems has made high-quality synthetic speech more accessible to the general public [4, 5].
@@
Although originally developed for positive applications, this technology is increasingly being exploited for malicious purposes [6, 7]. Synthetic speech generated with harmful intent, often referred to as spoofing, is being used to deceive individuals in scenarios such as voice phishing (or vishing). Spoofing also undermines the reliability of speech biometric systems, including Automatic Speaker Verification (ASV), many of which remain highly vulnerable to such attacks [8, 9].
@@
In response to these challenges, several datasets have been developed to advance research in Speech Deepfake Detection (SDD) [10, 11, 12]. For robust recognition systems, it is essential to have training data that cover a wide range of real-world acoustic environments and speaker diversity. However, speech generation systems, such as TTS and Voice Conversion (VC), typically require studio-quality or clean, read speech for training. Therefore, current datasets tend to feature clean, monotonic bona fide speech, with spoofed samples also being clean, as they are synthesized using TTS and VC systems trained on such data. The emerging task of Spoofing-robust Automatic Speaker Verification (SASV) [13] lacks dedicated datasets. Many SDD datasets also suffer from limited speaker diversity, which hinders research on SASV systems that require training with data from hundreds or even thousands of speakers.
@@
To this end, we introduce SpoofCeleb, a dataset built upon VoxCeleb1 [14], a widely used ASV dataset consisting of the voices of 1,25112511,2511 , 251 celebrities recorded under real-world conditions. We also develop a fully automated pipeline that processes VoxCeleb1 to produce in-the-wild bona fide speech samples that can be used for training TTS systems. [^1] The development of this pipeline is extensive, and the resulting bona fide speech data can serve other purposes, such as advancing research on TTS systems trained on noisy, in-the-wild data. We detail this aspect in a separate work, referring to the dataset as TTS In The Wild (TITW) [15]. From the two available TTS training sets in TITW, we use TITW-Easy as the source dataset to generate 23 spoofing attacks. SpoofCeleb is the first dataset explicitly designed for both SDD and SASV, where the bona fide speech is real-world, noisy speech. The dataset is divided into three subsets for training, validation, and evaluation, accompanied by evaluation protocols. Baseline systems trained on SpoofCeleb’s training set are also presented, demonstrating SpoofCeleb’s effectiveness in and potential for future research in SDD and SASV.
@@
Table 1: List of datasets in Speech Deepfake Detection (SDD) and Spoofing-robust Automatic Speaker Verification (SASV). FakeAVCeleb [16] and “In The Wild [17]” also have in-the-wild data. However, they have either only an evaluation set or the number of speakers or spoofing attacks is limited.
@@ 2 RELATED WORKS
Datasets for SDD and the generation-recognition trade-off. To safeguard the authenticity of speech, several datasets have been published to support research in SDD [10, 11, 18, 9, 26, 21, 12, 23]. One of the most critical decisions when creating these datasets is the selection of the source data (i.e., bona fide speech). This decision involves a trade-off, which we refer to as the “generation-recognition trade-off.”
@@
For both SDD and SASV on the recognition side, incorporating data with diverse noise, reverberation, and varied domains is essential for training robust models. It is well known that recognition models trained solely on clean speech often struggle to effectively generalize to noisy environments during inference [18]. While data augmentation techniques can help mitigate this issue [31], the most effective solution is to use training data drawn from a wide range of real-world sources.
@@
Conversely, traditional TTS training requires a carefully curated and recorded dataset. Sentence prompts must be selected to ensure comprehensive phonetic coverage [32], and recordings are typically made by voice professionals in clean environments, ideally in a single anechoic studio. These recordings are of high studio quality and carefully articulated but are not scalable. For instance, the well-known CMU Arctic database includes recordings from fewer than 10 voice professionals, each reading approximately 1,000 speech prompts [32]. Modern TTS systems, however, often require significantly more training data. Instead of relying on these small-scale, TTS-specific databases, contemporary models frequently use audiobook datasets (e.g., MLS [33]), which, while not studio-grade, consist of relatively clean audiobook recordings made by numerous readers in their homes or offices.
@@
Current SDD datasets tend to lean towards the generation side of the generation-recognition trade-off. They use source datasets that consist of either studio-quality or high-quality speech, facilitating the training of TTS and VC systems and the successful generation of spoofed speech samples. However, both the bona fide and spoofed speech in these datasets are exceedingly clean, making them far from real-world, noisy speech data.
@@
SpoofCeleb is the first dataset to use real-world, noisy, and reverberant data originating from TITW, which originates from VoxCeleb1, as the source for training and synthesizing spoofed speech. We tackle the generation-recognition trade-off by using our carefully curated, fully automated pre-processing pipeline that enables TTS models to be trained on data that more closely mirrors real-world conditions.
@@
Datasets for SASV. As SASV is an emerging task extending the scope of ASV systems with spoofing robustness, there is a lack of dedicated datasets for SASV. Earlier studies on SASV have relied on SDD datasets [34, 35]. However, current SDD datasets do not prioritize speaker diversity and balance, both of which are critical for SASV. Most datasets also lack a sufficient number of speakers.
@@
To the best of our knowledge, VSASV [30], a parallel data collection effort to SpoofCeleb, is the only attempt at addressing these limitations by creating a dataset specifically for SASV. SpoofCeleb complements VSASV while also having several distinctions. While VSASV includes three spoofing attacks, SpoofCeleb contains 23. Although VSASV uses in-the-wild bona fide data, its spoofed data are derived from high-quality sources due to the challenges in developing TTS systems with in-the-wild data. In contrast, SpoofCeleb adopts TITW which originates from VoxCeleb1, a widely-used ASV dataset recorded in the wild, as its bona fide source. Additionally, VSASV includes approximately 300 k samples, whereas SpoofCeleb offers over 2.5 M samples. Table 1 compares SpoofCeleb with other SDD and SASV datasets.
@@
Figure 1: Overall process pipeline of SpoofCeleb dataset collection. (a): our proposed fully automated pipeline transcribes, segments, filters, enhances, and again filters with DNSMOS to derive TITW-Easy [15] from VoxCeleb1 [14], which is adequate for TTS training. (b): 23 different TTS systems are trained using TITW-Easy and spoof speech samples are generated. All generated spoofing samples are combined with TITW-Easy to constitute SpoofCeleb.
@@ 3 SOURCE DATASET: TITW
Our goal is to create a dataset for SDD and SASV using VoxCeleb1 as the source so that both bona fide and spoofed samples would reflect real-world scenarios. However, VoxCeleb1 is not suitable for direct use in TTS training. [^2] Our preliminary attempts to train TTS systems using the raw VoxCeleb1 data without further processing were unsuccessful. The challenges with VoxCeleb1 are multifaceted. For example, the speech samples often (i) contain overly emotional expressions, (ii) include extended non-speech segments, or (iii) have excessively long durations. To address these issues, our developed fully automated pipeline processes VoxCeleb1 into the TITW dataset, which can be used for TTS training.
Figure 1-(a) illustrates the automated processing pipeline that was used to generate the TITW dataset. The pipeline begins by transcribing and obtaining word-level alignment using the WhisperX toolkit [36]. This toolkit transcribes the speech using the pre-trained Whisper Large v2 Automatic Speech Recognition (ASR) model [37], while word-level segmentation is derived from another phoneme-based ASR model. For a small subset of randomly selected samples, we also transcribe the text using the OWSMv3.1 model [38] and cross-check the accuracy of the transcriptions. We then segment the utterances from VoxCeleb1 whenever a silence longer than 500 ms is detected, resulting in multiple segments from a single utterance. Next, we apply a series of heuristic-driven rules – developed through several iterations of TTS training – to filter the data. We discarded any samples that (i) were non-English, (ii) were shorter than 1 second or longer than 8 seconds, (iii) contained one or more words with a duration exceeding 500ms, or (iv) had empty transcriptions.
After completing the initial processing steps (referred to as TITW-Hard in [15]), we conducted multiple iterations of TTS training trials. Despite these efforts, training remained extremely challenging for most TTS systems, with only a few recent models showing success. The generated speech was still insufficient to deceive pre-trained ASV systems, as measured using the SPooF Equal Error Rate (SPF-EER) metric [13]. [^3] The SPF-EER is calculated by assessing an ASV system’s ability to correctly accept target trials while rejecting spoofed non-target trials. Bona fide non-target trials are excluded from this protocol, as the focus is solely on evaluating the ASV system’s spoofing robustness. To address this, we applied speech enhancement using a pre-trained model named DEMUCS and excluded samples with DNSMOS “BAK” (background noisy quality) scores below 3.0. The final number of speech segments (TITW-Easy in [15]) is approximately 248 k, which serves as the bona fide portion of the SpoofCeleb dataset. For full details on the preparation of TITW from VoxCeleb1, refer to [15]. Nonetheless, we note that this choice of enhancing the bona fide speech may confuse the training of detection models because inevitable artifacts can be added with the enhancement process. Yet, we employ TITW-Easy as the bona fide, not TITW-Hard, because of the aforementioned practicality.
@@ 4 SPOOFCELEB
Figure 1-(b) illustrates the composition of SpoofCeleb. The TITW dataset serves as the foundation for training multiple TTS systems. These systems are then used to synthesize spoofed speech samples, which are combined with the bona fide speech samples from TITW to form the complete SpoofCeleb. To achieve this, we use 4 acoustic models, 6 waveform models (i.e., vocoders), and 5 End-to-End (E2E) models. Unless mentioned otherwise, all models were trained from scratch using the TITW-Easy data. SpoofCeleb does not include voice conversion systems, as TTS systems pose more immediate and prevalent security threats with publicly available APIs. Incorporating voice conversion systems would also require more complex configurations, such as defining source and target speaker pairs. Hence we leave this part for future work.
@@ 4.1 Acoustic models
Training acoustic models using in-the-wild data was one of the most challenging aspects of SpoofCeleb creation. We applied several criteria to evaluate the success of the training, including (but not limited to) speech intelligibility, measured by the Word Error Rate (WER), noisiness, assessed using DNSMOS, and speaker identity, evaluated using SPF-EER. Among these metrics, SPF-EER was prioritized as the primary measure, since the most critical factor in a spoofing attack is whether it can deceive an ASV system. The final models that were successfully trained include TransformerTTS, GradTTS, Matcha-TTS, and BVAE-TTS.
TransformerTTS  [39] is an autoregressive TTS model that generates mel-spectrograms from textual input using a transformer-based architecture. The model uses a sequence of transformer encoder and decoder blocks with multi-head self-attention. We trained TransformerTTS using the ESPnet toolkit [40]. [^4] https://github.com/ESPnet/ESPnet.
GradTTS.  [41] is a TTS model with a score-based decoder that generates mel-spectrograms by gradually transforming noise predicted by the text encoder. During inference, we set the denoise step to 50 to ensure high-quality speech generation. We used the official implementation and followed the default settings. [^5] https://github.com/huawei-noah/Speech-Backbones.
Matcha-TTS. [42] is an efficient non-autoregressive TTS model based on an optimal-transport conditional flow matching decoder [1]. Unlike score-based models, it constructs a more direct sampling trajectory, enabling high-quality generation with fewer sampling steps. We used the official implementation. [^6] https://github.com/shivammehta25/Matcha-TTS.
BVAE-TTS.  [43] uses a Bidirectional-inference Variational AutoEncoder (BVAE) to model the hierarchical relationships between text and speech. By leveraging the attention maps generated using BVAE-TTS, the model jointly trains a duration predictor, enabling robust and efficient non-autoregressive speech generation. We used the official implementation. [^7] https://github.com/LEEYOONHYUNG/BVAE-TTS.